OpenStudy (anonymous):

Write an equation for the circle that satisfies the given set of conditions. endpoints of a diameter at (11, 18) and (-13, -19)

OpenStudy (anonymous):

the whole question like I don't even know.

OpenStudy (anonymous):

The midpoint of the diameter is the center of the radius. Half the length of the diameter is the length of the radius. The equation of a circle with center at (h, k) and radius r is: \((x - h)^2 + (y - k)^2 = r^2\)

OpenStudy (anonymous):

no. :(

OpenStudy (anonymous):

if you show me how to do this one i can do the rest of my homework. lol

OpenStudy (anonymous):

The midpoint of a segment with endpoints \((x_1, y_1)\) and \((x_2 + y_2)\) is \(\left( \dfrac{x_1 + x_2}{2}, \dfrac{y_1 + y_2}{2} \right)\)

OpenStudy (anonymous):

The distance, d, between points \((x_1, y_1)\) and \((x_2, y_2)\) is \(d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\)

OpenStudy (anonymous):



OpenStudy (anonymous):

The figure above more or less represents this problem.

OpenStudy (anonymous):

@maddie1313 Are you following?

OpenStudy (anonymous):

yes. But how do I get an equation from the figure?